OpenStudy (anonymous):

How to find 3 sin 45 - 4 tan pi/6

OpenStudy (anonymous):

Find what sin 45 and tan(pi/6) is. Use a calculator if you don't have it memorized. sin 45=\[\frac{ 1 }{ \sqrt{2} }\] and tan(pi/6)=\[\frac{ 1 }{ \sqrt{3} }\]